How much energy is needed to heat 523 g of water from 20.0

o

oC to 92.5oC?  



1
Expert's answer
2021-03-04T06:35:16-0500

First of all, the melting process of the ice into water at 20 degree Celsius involves 2 phases. The 0 C ice firsts melts into 0 C water. Then following up is the phase of increasing temperature of the water from 0 C to 20 C.

The total amount of energy for the whole process is the sum of the energy used to change the state of the ice to water and the energy used to change the temperature from 0 C to 20 C. Writing mathematically,

Q=mL+mcθ

Q=mL+mcθ

where,

Q= the total amount of energy required throughout the process ( in the unit Joule, J)

m= mass of the ice cube( in the unit kilogram, kg)

L= the specific latent heat of fusion of ice (=3.36×10

5

3.36×105J kg

−

1

−1)

c= the specific heat capacity of water (= 4181 J kg

−

1

−1C 

−

1)

−1)

θ

θ= the change in temperature of the liquid (in the unit degree Celsius, C)

From the equation, we can calculate the total energy needed.

Q=0.01(3.36×10

5

)+0.01(4181)(20−0)

Q=0.01(3.36×105)+0.01(4181)(20−0)

Q=3360+8362

Q=3360+8362

Q=11722

Q=11722J

Therefore, the total amount of energy needed is 11722 Joules.